OpenStudy (anonymous):

How many solutions does the equation 5x + 10 = 5x − 8 have? Zero One Two Infinitely many

TheSmartOne (thesmartone):

Try to solve for x. If you get, just throwing random numbers: \(\sf 5 = 5\) Then you have infinitely many solutions If you get \(\sf x=3\) You have one solution. There can never be 2 solutions. If you get \(\sf 0 = -2\) since that isn't equal, you have 0 solutions.

OpenStudy (kitkat16):

to start of you need to change the x to 0 in the equation. 5(0)+10=5(0)-8 then multiply by 5 0+10=5(0)-8 do you follow so far?

OpenStudy (anonymous):

I think it's D...

OpenStudy (kitkat16):

I think so too! :)

OpenStudy (anonymous):

really? oh, well thank you lovely :)

TheSmartOne (thesmartone):

D is correct :)

OpenStudy (kitkat16):

because there are 3 possible solutions 0, 1, 2,

OpenStudy (anonymous):

thank you guys, i appreciate it :)

TheSmartOne (thesmartone):

(:

OpenStudy (kitkat16):

np :)